A 132mL of gas is measured at 38°C. If the pressure remains constant, what will be the volume of the gas at 10°C?


1
Expert's answer
2022-03-03T07:27:01-0500

At constant pressure,

V1÷T1 = V2÷T2

V1=132ml, T1= 38+273=311K

T2= 10+273= 283K


132÷311= V2÷283


V2=132×283÷311

=120.12 ml